OpenStudy (anonymous):

Someone HELP ME PLEEEASEEE!! what is the simplified form of the exression [2/5x^3]^2

OpenStudy (anonymous):

\[\left( \frac{ 2 }{ 5x^3 } \right)^2 \implies \frac{ 2^2 }{ (5x^3)^2 }\]

OpenStudy (anonymous):

is that the final answer ? @iambatman

OpenStudy (anonymous):

Nope, you have to simplify it more, that's why I left it for you

OpenStudy (anonymous):

Remember \[(x^n)^m \implies x^{nm}\]

OpenStudy (anonymous):

so its gonna be 2^2/5x^22 ? @iambatman

OpenStudy (anonymous):

No, what is 2^2?

OpenStudy (anonymous):



OpenStudy (anonymous):

Yes, what does that equal?

OpenStudy (anonymous):

Do you know what that means

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Yes, so now what's (5x^3)^2

OpenStudy (anonymous):

125x^2 ?

OpenStudy (anonymous):

i think i got the wrong answer

OpenStudy (anonymous):

\[(5x^3)^2 \implies 25x^6\]

OpenStudy (anonymous):

Do you see how I got that?

OpenStudy (anonymous):

how did u get 25 ?

OpenStudy (anonymous):

\[(5x^3)^2 \implies 5^2x^{3 \times 2} \implies 25x^6\]

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

oh okay, so then will we multiply 25 by 6 or is that the final answer

OpenStudy (anonymous):

No, you don't multiply by the variable, that x does not mean multiply, it's a variable, like it could be, y, z, b, what ever.

OpenStudy (anonymous):

\[\frac{ 4 }{ 25x^6 }\] this would be your answer, but next time don't worry about the answer too much, try to figure out how to simplify the best you can.

OpenStudy (anonymous):

Knowing the answer doesn't help, it's doing the problem itself that will help you learn.

OpenStudy (anonymous):

so we have to follow the PEDMAS rule right ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

But that's your final answer

OpenStudy (anonymous):

Unless you want decimals

OpenStudy (anonymous):

okay so theres this other problem [t^4/7y^3]^-3 how do i do it with letters ?

OpenStudy (anonymous):